Which of the following numbers does not fit into the series?
14, 19, 29, 40, 44, 51, 59, 73

Grade:12th pass

1 Answers

Arun
4 years ago

1+4= 5 , 14 + 5 = 19

1+9= 10 , 19 + 19 = 29

2+9= 11 , 29 + 11 = 40

40 + 12 = 52

hence 51 is incorrect